# Estimating the linear relation between variables that are never jointly observed: an application in in vivo experiments

### v1

@@ 3 Bootstrapping for confidence intervals
Within each group $k$, we draw, with equal probability and with replacement, $n_{y}^{k}$ observations among $Y_{1}^{k},\ldots,Y_{n_{y}^{k}}^{k}$, We also draw independently, with equal probability and with replacement, $n_{x}^{k}$ observations among $\mathbf{X}_{1}^{k},\ldots,\mathbf{X}_{n_{x}^{k}}^{k}$. We denote by $\mu_{Y}^{k*}$ and by $\boldsymbol{\mu}_{X}^{k*}$ the empirical means and by $\sigma_{Y,k}^{2,*}$ and by $\boldsymbol{\Gamma}_{X}^{k*}$ the empirical variances of $Y$ and $\mathbf{X}$ in these these bootstrap samples. Bootstrapped estimators $\boldsymbol{\beta}^{M,*}$ and $\boldsymbol{\beta}^{W,*}$ of $\boldsymbol{\beta}$ can now be computed by replacing the empirical moments by the bootstrap moments in (4) and (6).
@@
Suppose that the assumptions of Proposition 2.2 as fulfilled. As $n\to+\infty$, the bootstrap estimator $(\boldsymbol{\beta}^{W,*},\sigma_{\epsilon}^{2,W,*})$ is strongly consistent for $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ in the Kolmogorov metric. Consider a risk $\alpha\in(0,1)$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides, for a given nominal level $1-\alpha$, a consistent confidence set.
Propositions 3.1 and 3.2 are very important for practical applications since they ensure that even if we are not able to compute the explicit expression of the asymptotic variance of our estimators, we can still construct asymptotic confidence intervals thanks to simple bootstrap procedures.
4 A simulation study
Some simulations are performed in order to evaluate the finite sample performances of the proposed approaches on data that resemble in vivo data originating from real experiments on mice.
@@ 4.1 Simulation design
4.1.1 General setting
Let $K\in\mathbb{N}$ be the number of subpopulations, $g_{k}=\{0,1,\dots,k-1\}$ be the level of the grouping variable for subpopulation $k\in\{1,\dots,K\}$, and the number of animals observed per group was choose for simplicity $n=n_{y}^{1}=\cdots=n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$. For every animal $i\in\{1,\dots,n\}$ and every subpopulation $k\in\{1,\dots,K\}$ we choose the predictor variable $X^{k}_{i}$ in the Gaussian univariate setting:
$X^{k}_{i}\sim\mathcal{N}(\mu_{X}^{k},\sigma_{X,k}^{2})\text{, where }\mu_{X}^{k}=\Delta_{\mu}g_{k}+C_{\mu},$
where $C_{\mu}\in\mathbb{R}$ is the global mean scale parameter, and $\Delta_{\mu}\in\mathbb{R}$ is the parameter that determines the difference between groups in terms of the mean. We simulate the predicted variable independently as follows, with regression parameters $\beta_{0}$ and $\beta_{1}$:
$Y^{k}_{i}=\beta_{0}+\beta_{1}X^{\prime k}_{i}+\epsilon^{k}_{i}\text{, where }X^{\prime k}_{i}\sim X^{k}_{i}\text{ and }\epsilon^{k}_{i}\sim\mathcal{N}(0,\sigma_{\epsilon,k}^{2}).$
It should be noted that $X^{k}_{i}$ and $X^{\prime k}_{i}$ are independent, we thus recreate the situation of the predictor and predicted variables not being simultaneously observed.
The variable sets $X=(X^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ and $Y=(Y^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ are simulated $N_{sim}\in\mathbb{N}$ times. For each simulation, $N_{boot}\in\mathbb{N}$ bootstrap samples of the size $n$ are generated from $X^{k}=(X^{k}_{i})_{1\leq k\leq K}$ and $Y^{k}=(Y^{k}_{i})_{1\leq k\leq K}$ for each subpopulation $k\in\{1,\dots,K\}$ independently, then moment estimators $\mu_{X}^{k*}$ and $\mu_{Y}^{k*}$ are calculated. Finally, the bootstrap sample-based estimators $\boldsymbol{\beta}^{M,*}=(\beta_{0}^{M,*},\beta_{1}^{M,*})$ and $\boldsymbol{\beta}^{W,*}=(\beta_{0}^{W,*},\beta_{1}^{W,*})$ are calculated. Based on $N_{boot}$ estimates, we calculate 95% confidence intervals using the function quantile of the Python library NumPy. As a result, we obtain $N_{sim}$ confidence intervals for each regression parameter, that we use to calculate the following quantities of interest: coverage rate of the intervals, their average amplitude, and power, i.e. the proportion of intervals not including 0. Along with the bootstrap procedure, we estimate the confidence intervals for the parameter estimators of the regression on the means by group with a naive method, assuming that the deviation of the parameter estimator from the true value divided by the estimator’s standard error follows a Student’s t-distribution:
@@ 4.1.1 General setting
4.1.2 Parameter levels
We fix the following parameters throughout all simulations: the regression parameters $\beta_{0}=1$ and $\beta_{1}=2$, the number of simulations $N_{sim}=500$, the number of bootstrap samples $N_{boot}=500$, the location parameters of the predictor variable $\Delta_{\mu}=1$ and $C_{\mu}=10$. Multiple parameters are varied to study their effect. We take the number of animals $n\in\{10,30\}$, in particular to test whether the inference is significantly impaired in case of a small number of animals, which is often the case for real experimental data. We consider the number of groups $K\in\{4,10\}$, 4 being the number of groups that is often observed in real data, and 10 being a higher number that may produce sufficiently good results with the naive approach to approximating confidence intervals with Student distribution. Additionally, the parameter $\sigma^{2}_{X,k}$ can be adjusted to control the extent to which the observations per group can be easily distinguish one from another. We set $\sigma_{X,1}^{2}=\dots=\sigma_{X,K}^{2}$, and $\sigma^{2}_{X,k}\in\{0.75,2\}$, the first value corresponding to lower overlap between groups, and the second to higher overlap. Finally, we introduce an additional parameter $\rho\in\mathbb{R}^{+}$ controlling the variance of the response to the variance of noise ratio, i.e. $\rho=\cfrac{\sigma_{Y,k}}{\sigma_{\epsilon,k}}$. The choice of adjusting the noise to signal ratio instead of the quantity of noise itself through $\sigma_{\epsilon,k}^{2}$ is motivated by the fact that $\sigma_{Y,k}$ depends on $\sigma_{X,k}$, hence the same level of $\sigma_{\epsilon,k}$ cannot be interpreted the same way for different values of $\sigma_{X,k}$. The variance of the noise can be expressed as follows: $\sigma_{\epsilon,k}^{2}=\cfrac{\beta_{1}^{2}\sigma_{X,k}^{2}}{\rho^{2}-1}$. The values of $\rho$ are chosen to correspond to the realistic situation, namely the very noisy case and a slightly less noisy one: $\rho\in\{1.01,1.1\}$. The effect of different values of $\rho$ on the simulated response variable is illustrated on Figure 3.
Figure 3: The effect of different values of $\rho$ on the data, with $K=4$ and $\sigma_{X,1}^{2}=\dots=\sigma_{X,4}^{2}=0.75$. a) Boxplots constructed from the simulated values of $X^{k}_{i}$. b) Boxplots constructed from the simulated values of $Y^{k}_{i}$ with lower relative noise level, i.e. $\rho=1.1$. c) Boxplots constructed from the simulated values of $Y^{k}_{i}$ with higher relative noise level, i.e. $\rho=1.01$.
@@ 4.2 Results
The results of the simulation study are presented in Figure 4, namely the obtained coverage rates for the linear regression slope confidence intervals on Figure 3(a), the amplitudes on Figure 3(b), and the associated test powers on Figure 3(c). Every figure contains four tables, each corresponding to a combination of parameters with respect to the number of animals $N_{anim}$ and the number of groups $K$. The columns of the tables represent combinations of parameters with respect to the group dispersion parameter $\sigma^{2}_{X,k}$ and the noise to signal ratio $\rho$. Finally, the lines of each table indicate the method used to estimate the confidence intervals: the proposed method of moments and optimal transport based bootstrap estimators, and the classical linear regression procedure based on estimated means within each group.
Figure 4: a) Coverage rates, b) amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from $N_{sim}$ simulations for each parameter combination. The method labels on the left: "mm (boot)" stands for the method of moments with bootstrap, "ot (boot)" for the optimal transport method with bootstrap, and "mm (student)" for the naive linear regression on means approach based on Student’s distribution.
@@
In general, it can be observed that the bootstrap estimators produce confidence intervals with smaller amplitudes and with higher power, but with lower coverage rate. Whereas the coverage rates in all cases remain within 91-96%, the extent to which the amplitudes are smaller and the powers are bigger is significantly more important for the bootstrap estimators in almost all cases. This implies that the naive approach based on the Student’s distribution is more likely to produce false negatives in terms of significance.This trend of further amplified by the number of groups parameter: whereas the results are overall worsened with the decrease in either the number of animals or groups, it is the case with the small number of groups that demonstrates the biggest difference in the approaches. Indeed, in almost all cases within the tables with $K=4$ we observe the amplitudes approximately twice as important for the naive approach, and a similar trend in terms of lower powers. The latter result is important since lower power implies bigger probability of not detecting a significant relationship between the predictor and the predicted variables, when it is actually present. Overall, these results mean that the proposed bootstrap estimators are more effective if the experimental design entails a small number of groups.
@@
Concerning the remaining two parameters, as expected, in general the best results are obtained with lower $\sigma^{2}_{X,k}$ and higher $\rho$. In most case, the results for the naive and the bootstrap estimators are either both good or both bad in terms of power, with the latter being slightly better. A particularly complicated case can be distinguished, with high dispersion and high noise level, with few groups and few animals, where all estimators drastically fail: we observe almost equally bad powers (0.1 for both bootstrap estimators and 0.09 for the naive estimator), despite the significant difference in amplitudes. On the other hand, we can also distinguish two case where the powers of the bootstrap estimators are more than 90%, whereas those of the naive estimator are under 50%: in both cases there are 4 groups and high noise, in the first case there are only 10 animals but lower dispersion, in the second case high dispersion level is compensated but a higher number of animals. This implies that if the underlying distributions per group are characterized by a reasonable amount of overlap, or a significant overlap is compensated by having more observations, the bootstrap estimators manage to detect the significant relationship in most cases, unlike the naive estimator.
@@
Lastly, it can be observed that the estimator based on optimal transport produces confidence intervals with slightly smaller amplitudes compared to the method of moments estimator. The difference appears to be relatively more important in the cases with higher group overlap $\sigma^{2}_{X,k}=2$. However, the powers are not affected by this difference. This may be explained by a more important bias associated with the optimal transport estimator. The estimator will likely produce better results in terms of the power than the method of moments estimator if the bias is corrected.
@@ 5 Application to real data
In order to illustrate the proposed estimators on real data, we studied the data from experiments conducted on mice in order to assess the adverse effects induced in the context of different irradiated volume. In the course of these experiments, mice were exposed to either stereotactic body radiation therapy (SBRT) with different beam sizes at 90 Gy on the left lung, or whole-thorax irradiation (WTI) at 19 Gy. For one cohort, the expressions of three pro-inflammatory genes (IL1 $\alpha$ and $\beta$, IL6 and TNF) were measured, for the other cohort the measurements of the thickness of the alveolar septas were made due to its role as a macroscopic biomarker of radio-induced pulmonary lesions. In the case of the SBRT, the measurements were taken in multiple locations: the irradiated patch (inside the irradiation field), the remaining part of the left lung referred to as ipsilateral lung, and the right lung (contralateral lung). The goal of this statistical analysis is to determine whether there is a statistical association between the gene expression as predictors and the septal thickening as as outcome. Our approach is applied since the variables are measured on different animals, but within each irradiation condition there are shared groups in terms of the measurement time points.
Table 1: Results of estimation of the linear regression slope predicting septal thickening with the pro-inflammatory genes expression, with three methods, for WTI and SBRT with different beam sizes, with measurements taken in different parts of lungs.
@@
IL1b -0.92 (-2.2, 1.06) -0.42 (-0.89, 0.51) -0.92 (-3.97, 2.14)
0 mm IL6 2.23 (-1.99, 2.28) 0.83 (-0.83, 0.97) 2.23 (-2.65, 7.12)
TNF 2 (-2.0, 2.82) 0.88 (-1.0, 1.11) 2 (-1.72, 5.72)
IL1b 0.63 (-0.82, 2.32) 0.23 (-0.33, 0.56) 0.63 (-2.95, 4.21)
1 mm IL6 1.03 (-0.57, 2.19) 0.46 (-0.3, 1.0) 1.03 (-0.88, 2.94)
Right TNF 1.05 (-0.47, 2.43) 0.66 (-0.31, 1.26) 1.05 (-1.01, 3.11)
lung IL1b 1.07 (0.17, 1.97) 0.92 (0.17, 1.13) 1.07 (-1.96, 4.1)
3 mm IL6 0.3 (-1.45, 1.12) 0.37 (-0.74, 0.86) 0.3 (-4.94, 5.53)
TNF 2.02 (0.27, 4.1) 1.05 (0.04, 1.44) 2.02 (0.03, 4.02)
IL1b 6.41 (-11.6, 15.4) 1.62 (-2.28, 3.1) 6.41 (-16.7, 29.52)
7 mm IL6 -0.58 (-1.26, 0.09) -0.6 (-1.23, 0.1) -0.58 (-6.23, 5.08)
TNF 0.51 (-1.7, 3.86) 0.43 (-1.45, 1.89) 0.51 (-23.21, 24.22)
Whole IL1a 2.93 (-1.18, 6.63) 2.47 (-1.49, 3.02) 2.93 (-11.08, 16.94)
thorax IL6 3.7 (1.53, 5.99) 2.51 (1.39, 3.43) 3.7 (1.11, 6.29)
@@
The linear regression parameters were estimated with three estimators the same way it was done in the simulation study in Section 4.2. The results are presented in Table 1. The focus is placed on estimating the slope parameter $\beta_{1}$ in particular. The table contains the estimations of $\beta_{1}$ as well as the estimated confidence intervals for the slope estimator, and the corresponding test result on the significance of the estimated relationship.
Figure 5: Distribution of the data, collected from the irradiated patch under SBRT with 7 mm beam size: the expression of the gene TNF on the left, and septal thickness on the right. The measurements were made 30, 45 and 60 days after irradiation.
@@
Major differences between the results for the naive and the bootstrap estimators can be observed in terms of the detected significance. On the one hand, the relationship between the pro-inflammatory genes and septal thickening has been identified by all methods in the case of whole-thorax irradiation (two out of three genes for the bootstrap estimators, and only one for the naive estimator), which is an expected result. On the other hand, we also expect to identify a strong correlation in the case of the measurements taken directly from the irradiated patch. This is only the case for the bootstrap estimators, but not for the naive one. This results is in accordance with the results we obtain with simulated data: the confidence intervals are often overestimated with the naive approach, which may result in false negatives in terms of significance.
@@
For the irradiated patch, as well as in all other cases with identified significance, it is only the case for the beam size of 3 mm, the results that is consistent with literature, indicating it as the beam size starting with which the long-term lesions start appearing (Bertho et al.,, 2020). Multiple significant associations have been identified with the bootstrap estimators in the ipsilateral lung and in the right lung for the beam size of 3 mm. However, none of the genes has been identified as significantly linked to lesions in the case of 7 mm. This may be caused by the relationship being of non-linear nature, as appears to be the case for the gene TNF, presented in Figure 5.
@@
Finally, among the cases where a significant relationship has been detected, the estimated values of the slope are always positive, which indicates a general radio-induced up-regulation trend. These values are in general bigger in case of the whole-thorax irradiation and within the patch than for the ipsilateral or right lung for SBRT, which is also in line with biological knowledge, since the genes are expected to be more strongly up-regulated in the areas of injury than further away. This effect is illustrated in Figure 6 on the example of the linear model prediction made for the gene IL6 .
@@
Figure 6: Linear model prediction of septal thickness based on IL6 expression, plotted for different locations and beam sizes, with the results from two bootstrap estimators.

### v3

@@ 6 Bootstrapping for confidence intervals
Within each group $k$, we draw, with equal probability and with replacement, $n_{y}^{k}$ observations among $Y_{1}^{k},\ldots,Y_{n_{y}^{k}}^{k}$, We also draw independently, with equal probability and with replacement, $n_{x}^{k}$ observations among $\mathbf{X}_{1}^{k},\ldots,\mathbf{X}_{n_{x}^{k}}^{k}$. We denote by $\mu_{Y}^{k*}$ and by $\boldsymbol{\mu}_{X}^{k*}$ the empirical means and by $\sigma_{Y,k}^{2,*}$ and by $\boldsymbol{\Gamma}_{X}^{k*}$ the empirical variances of $Y$ and $\mathbf{X}$ in these bootstrap samples. Bootstrapped estimators $\boldsymbol{\beta}^{M,*}$ and $\boldsymbol{\beta}^{W,*}$ of $\boldsymbol{\beta}$ can now be computed by replacing the empirical moments by the bootstrap moments in (3.2) and (3.4).
@@
Suppose that the assumptions of Proposition 4.2 are fulfilled. As $n\to+\infty$, the bootstrap estimator $(\boldsymbol{\beta}^{W,*},\sigma_{\epsilon}^{2,W,*})$ is strongly consistent for $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$ in the Kolmogorov metric. Consider a risk $\alpha\in(0,1)$, for each component of $\boldsymbol{\beta}$, the bootstrap percentile approach provides, for a given nominal level $1-\alpha$, a consistent confidence set.
Propositions 6.1 and 6.2 are very important for practical applications since they ensure that even if we are not able to compute the explicit expression of the asymptotic variance of our estimators, we can still construct asymptotic confidence intervals thanks to simple bootstrap procedures.
7 Simulation study
@@ 7.1 Simulation design
We performed a series of simulations in order to evaluate the finite sample performances of the proposed approaches on data that resemble in vivo data originating from real experiments on mice. The number of animals observed per group is chosen for simplicity $n=n_{y}^{1}=\cdots=n_{y}^{K}=n_{x}^{1}\cdots=n_{x}^{K}$, and thus the weights for every group are also considered to be equal, i.e. $w_{1}=\cdots=w_{K}=\frac{1}{K}$. For every animal $i\in\{1,\dots,n\}$ and every subpopulation $k\in\{1,\dots,K\}$, the predictor variable $X^{k}_{i}$ is Gaussian univariate: $X^{k}_{i}\sim\mathcal{N}(\mu_{X}^{k},\sigma_{X}^{2})$, where $\mu_{X}^{k}=9+k$. We simulate the predicted variable independently as $Y^{k}_{i}=\beta_{0}+\beta_{1}X^{\prime k}_{i}+\epsilon^{k}_{i}$, where $X^{\prime k}_{i}\sim X^{k}_{i}$ and $\epsilon^{k}_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, with regression parameters $\beta_{0}=1$ and $\beta_{1}=2$. It should be noted that $X^{k}_{i}$ and $X^{\prime k}_{i}$ are independent, we thus recreate the situation of the predictor and predicted variables not being simultaneously observed.
The variable sets $X=(X^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ and $Y=(Y^{k}_{i})_{1\leq i\leq n,1\leq k\leq K}$ are simulated $N_{sim}$ times. For each simulation, $N_{boot}$ bootstrap samples of the size $n$ are generated from $X^{k}=(X^{k}_{i})_{1\leq k\leq K}$ and $Y^{k}=(Y^{k}_{i})_{1\leq k\leq K}$ for each subpopulation $k\in\{1,\dots,K\}$ independently, then moment estimators $\mu_{X}^{k*}$ and $\mu_{Y}^{k*}$ are calculated. Finally, the bootstrap sample-based estimators $\boldsymbol{\beta}^{M,*}=(\beta_{0}^{M,*},\beta_{1}^{M,*})$ and $\boldsymbol{\beta}^{W,*}=(\beta_{0}^{W,*},\beta_{1}^{W,*})$ are calculated. Based on $N_{boot}$ estimates, we calculate 95% confidence intervals using the function quantile of the Python library NumPy. As a result, we obtain $N_{sim}$ confidence intervals for each regression parameter, that we use to calculate the following quantities of interest: coverage rate of the intervals, their average amplitude, and power, i.e. the proportion of intervals not including 0. Along with the bootstrap estimators, we also considered asymptotic confidence intervals in case of the method of moments, obtained by plugging the estimated values of the moments into the expression for the limit distribution presented in Lemma 5.1. Additionally, we estimate the confidence intervals for the parameter estimators of the regression on the means by group with a naive method, assuming that the deviation of the parameter estimator from the true value divided by the estimator’s standard error follows a Student’s t-distribution:
@@
Figure 7.1: The effect of different values of $\rho$ on the data, with $K=4$ and $\sigma_{X}^{2}=0.75$. a) Boxplots constructed from the simulated values of $X^{k}_{i}$. b) Boxplots constructed from the simulated values of $Y^{k}_{i}$ with lower relative noise level, i.e. $\rho=1.1$. c) Boxplots constructed from the simulated values of $Y^{k}_{i}$ with higher relative noise level, i.e. $\rho=1.01$.
Throughout all simulations, we fix the number of simulations $N_{sim}=500$ and the number of bootstrap samples $N_{boot}=500$. Multiple parameters are varied to study their effect. We take the number of animals $n\in\{10,30\}$, in particular to test whether the inference is significantly impaired if the measurements are available only for a small number of animals, which is often the case for real experimental data. We consider the number of groups $K\in\{4,10\}$, 4 being the number of groups that is often observed in real data, and 10 being a higher number that may produce sufficiently good results with the naive approach to approximating confidence intervals with Student distribution. Additionally, the parameter $\sigma^{2}_{X}$ can be adjusted to control the extent to which the observations per group can be easily distinguished one from another. We set $\sigma^{2}_{X}\in\{0.75,2\}$, the first value corresponding to lower overlap between groups, and the second to higher overlap. Finally, we introduce an additional parameter $\rho\in\mathbb{R}^{+}$ controlling the variance of the response to the variance of noise ratio, i.e. $\rho=\frac{\sigma_{Y}}{\sigma_{\epsilon}}$, where $\sigma^{2}_{Y}=Var(Y^{k}_{i})$ for all $i$ and $k$. The choice of adjusting the noise to signal ratio instead of the quantity of noise itself through $\sigma_{\epsilon}^{2}$ is motivated by the fact that $\sigma_{Y}$ depends on $\sigma_{X}$, hence the same level of $\sigma_{\epsilon}$ cannot be interpreted the same way for different values of $\sigma_{X}$. The variance of the noise can be expressed as follows: $\sigma_{\epsilon}^{2}=\frac{\beta_{1}^{2}\sigma_{X}^{2}}{\rho^{2}-1}$. The values of $\rho$ are chosen to correspond to the realistic situation, namely a very noisy case and a slightly less noisy one: $\rho\in\{1.01,1.1\}$. The effect of different values of $\rho$ on the simulated response variable is illustrated on Figure 7.1.
@@ 7.2 Results
The results of the simulation study are presented in Supplementary Figure C.1. Firstly, the results for the moment approach in the asymptotic and bootstrap cases are very similar, thus confirming the effectiveness of the bootstrap approach. Concerning the comparison between the bootstrap procedure and the naive approach, it can be observed that in general the bootstrap estimators produce confidence intervals with smaller amplitudes and with higher power, at the expense of a slightly lower coverage rate. Whereas the empirical coverage rates in all cases are close to the nominal one (95%), the extent to which the amplitudes are smaller and the powers are bigger for the bootstrap estimators is very important in almost all cases. This implies that the naive approach based on the Student’s distribution is more likely to produce false negatives in terms of significance. This trend is further amplified by the number of groups parameter: whereas the results are overall worsened with the decrease in either the number of animals or groups, it is the case with the small number of groups that demonstrates the biggest difference in the approaches (the case with few groups and animals is presented in Figure 7.2). Indeed, in almost all cases within the tables with $K=4$ we observe the amplitudes approximately twice as important for the naive approach, and a similar trend in terms of lower powers. The latter result is important since lower power implies bigger probability of not detecting a significant relationship between the predictor and the predicted variables, when it is actually present. Overall, these results mean that the proposed bootstrap estimators are more effective if the experimental design entails a small number of groups.
Figure 7.2: a) Coverage rates, b) amplitudes, and c) powers of the confidence intervals for the estimators of $\beta_{1}$ obtained from 500 simulations, with number of groups $K=4$ and number of animals per group $n=10$. The columns of the tables represent combinations of parameters with respect to the group dispersion parameter $\sigma^{2}_{X}$ and the noise to signal ratio $\rho$. The lines indicate the method used to estimate the confidence intervals: "mm (asymp)" stands for the method of moments with asymptotic confidence intervals, "mm (boot)" for the method of moments with bootstrap, "ot (boot)" for the optimal transport method with bootstrap, and "mm (student)" for the naive linear regression on means approach based on Student’s distribution.
@@
Concerning the remaining two parameters, as expected, in general the best results are obtained with lower $\sigma^{2}_{X}$ and higher $\rho$. In most cases, the results for the naive and the bootstrap estimators are either both good or both bad in terms of power, with the latter being slightly better. A particularly complicated case can be distinguished, with high dispersion, high noise level, few groups and few animals, where all estimators drastically fail: we observe almost equally bad powers (0.1 for both bootstrap estimators and 0.09 for the naive estimator), despite the significant difference in amplitudes. On the other hand, we can also distinguish two cases where the powers of the bootstrap estimators are more than 90%, whereas those of the naive estimator are under 50%: in both cases there are 4 groups and high noise, in the first case there are only 10 animals but lower dispersion, in the second case high dispersion level is compensated but a higher number of animals. This implies that if the underlying distributions per group are characterized by a reasonable amount of overlap, or a significant overlap is compensated by having more observations, the bootstrap estimators manage to detect the significant relationship in most cases, unlike the naive estimator.
@@
Lastly, it can be observed that the estimator based on optimal transport produces confidence intervals with slightly smaller amplitudes compared to the method of moments estimator. The difference appears to be relatively more important in the cases with higher group overlap $\sigma^{2}_{X}=2$. However, the powers are not affected by this difference. This may be explained by a more important bias associated with the optimal transport estimator. The estimator will likely produce better results in terms of the power than the method of moments estimator if the bias is corrected.
@@ 8 Application to real data
In order to illustrate the proposed estimators on real data, we studied the data previously mentioned in Section 1, obtained from experiments conducted on mice in order to assess the adverse effects induced in the context of different irradiated volume. In the course of these experiments, mice were exposed to either stereotactic body radiation therapy (SBRT) with different beam sizes at 90 Gy on the left lung, or whole-thorax irradiation (WTI) at 19 Gy. For one mouse cohort, the expressions of tho pro-inflammatory genes (IL6 and TNF) were measured, for the other cohort the measurements of the thickness of the alveolar septas were made due to its role as a severity measure of radio-induced pulmonary lesions. In the case of the SBRT, the measurements were taken in multiple locations: the irradiated patch (inside the irradiation field), the remaining part of the left lung referred to as ipsilateral lung, and the right lung (contralateral lung). A control condition is also considered, where the gene expressions are measured without prior radiation exposure. The goal of this statistical analysis is to determine whether there is a statistical association between the gene expression as a predictor and the septal thickening as as outcome. Our approach is applied since the variables are measured on different animals, but within each irradiation condition there are shared groups in terms of the measurement time points.
Table 8.1: Results of estimation of the linear regression slope predicting septal thickening with the pro-inflammatory genes expression, with three methods, for control (no irradiation), WTI and SBRT with different beam sizes, with measurements taken in different parts of lungs.
@@
Whole thorax IL6 3.7 (1.53, 5.99) 2.51 (1.39, 3.43) 3.7 (1.11, 6.29)
@@
In order to ensure the comparability of the results for different treatment conditions and genes, the data were centered and reduced with respect to the global mean and standard deviation prior to estimation. The linear regression parameters were then estimated with three estimators the same way it was done in the simulation study in Section 7. The focus is placed on estimating the slope parameter $\beta_{1}$ in particular. The results are presented in Table LABEL:tab:vivo_results, containing the estimations of $\beta_{1}$ as well as the estimated confidence intervals for the slope estimator, and the corresponding test result on the significance of the estimated relationship.
@@
Major differences between the results for the naive and the bootstrap estimators can be observed in terms of the detected significance. On the one hand, the relationship between the pro-inflammatory genes and septal thickening has been identified by all methods in the case of whole-thorax irradiation (both genes for the bootstrap estimators, and only one for the naive estimator), which is an expected result. The results are also consistent for all methods in the case without radiation exposure (control), where no significant relationship has been identified, as it is expected. On the other hand, we expect to identify a strong correlation in the case of the measurements taken directly from the irradiated patch. This is only the case for the bootstrap estimators, but not for the naive one. This results is in accordance with the results we obtain with simulated data: the confidence intervals are often overestimated with the naive approach, which may result in false negatives in terms of significance.
@@
Among the SBRT irradiation configurations, only the 3mm beam size has shown a significant correlation between inflammatory genes and septal thickening. These results are consistent with literature, indicating it as the beam size starting with which the long-term lesions appear (Bertho et al.,, 2020). Multiple significant associations have been identified with the bootstrap estimators in the ipsilateral lung and in the right lung for the beam size of 3 mm.
@@
Finally, among the cases where a significant relationship has been detected, the estimated values of the slope are always positive, which indicates a general radio-induced up-regulation trend. These values are in general bigger in case of whole-thorax irradiaion and within the patch than for the ipsilateral or right lung for SBRT, which suggests a stronger correlation between the inflammatory process and lung injuries under high dose/volume irradiation conditions. These results, which are in line with the biological knowledge, could not have been established within the framework of classical statistical regression approaches due to non-simultaneous observations. This effect is illustrated in Figure 8.1 on the example of the linear model prediction made for the gene IL6.
@@
Figure 8.1: Linear model prediction of septal thickness based on IL6 expression, plotted for different locations and beam sizes, with the results from two bootstrap estimators.